OpenStudy (anonymous):

Hi :) I need to resolve equation f(x)=0 in the intervale (-pi,pi) f(x)=sin(2x-pi/2)-1/2 I came here but i am not sure if it´s right. sin(2x)-cos(2x)-3/2=0 Can anyone help, please ? :)

OpenStudy (anonymous):

i help you

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

what grade you are in

OpenStudy (anonymous):

plesae

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what are are you in

OpenStudy (anonymous):

?

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

i am in second but i am not sure if it s same as in your country cause i am from Slovakia. I am 16 years old if it helps.

OpenStudy (anonymous):

your in second grade

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

why theying giving you that kind of work

OpenStudy (anonymous):

idk :) we have diffrent system (i am in bilingual class)

OpenStudy (freckles):

\[\sin(2x-\frac{\pi}{2})-\frac{1}{2}=0 \\ \sin(2x-\frac{\pi}{2})=\frac{1}{2} \\ \text{ Let } u=2x-\frac{\pi}{2} \\ \text{ so since we had } -\pi<x<\pi \text{ and } x=\frac{u}{2}+\frac{\pi}{4} \\ \text{ Then we need \to solve } \sin(u)=0 \text{ \in the interval } -\pi<\frac{u}{2}+\frac{\pi}{4}<\pi \\ -4\pi<2u+\pi<4\pi \\ -5\pi<2u<3\pi \\ \frac{-5\pi}{2}<u<\frac{3\pi}{2}\]

OpenStudy (freckles):

So you need to solve sin(u)=0 in the interval (-5pi/2,3pi/2)

OpenStudy (anonymous):

-.-

OpenStudy (freckles):

This is one way to solve your equation.

OpenStudy (anonymous):

i wont even understand that

OpenStudy (anonymous):

http://www.tutorpace.com/guest/classroom/4463

OpenStudy (anonymous):

thank you :)

OpenStudy (freckles):

hope you seen the type-o above solve sin(u)=1/2 in the interval (-5pi/2,3pi/2) *

OpenStudy (anonymous):

yes thank you really much :) it helped a lot :)